# Memristor-Based MobileNetV3 Circuit Design for Image Classification

### v2

@@ III Experiments
We refer to methods previously reported in the papers to analyze the latency and energy consumption of memristor-based neural networks [12, 15, 19, 20], to determine the latency of our proposed computing paradigm. The response time can be as quick as 100 ps [15]. Therefore, the latency of this circuit can be as low as 1.24 $\mu$s. This is significantly faster than the traditional GPU approach, which has a latency of 165.4 $\mu$s on a server equipped with an RTX 4090. Figure 6 illustrates the distribution of memristor weights across different layers. As observed in Figure 6, the weights of the memristor crossbars predominantly range between -0.2 and 0.2. The power consumption of the memristor-based neural network is derived from the weights of the memristors. Since we have mapped the input data to ±2.5 mV, we can estimate the maximum predicted power consumption of a single memristor by assuming that all input data is at 2.5 mV and all weights are 0.2. Therefore, the maximum power consumption of memristors can be estimated at 1.1 $\mu$W, and the power consumption of the CMOS circuit required to achieve the same calculation is close to 60 $\mu$[21].
@@
IV Conclusion
The memristor-based MobileNetV3 circuit design proposed in this paper demonstrates exceptional performance in accuracy, latency, and energy efficiency for image classification tasks on edge devices. Our optimized memristive modules leverage the advantages of memristors to achieve over 90% accuracy on the CIFAR-10 dataset while significantly reducing inference time and power consumption compared to traditional implementations.

### v3

@@ IV-A Accuracy
In this section, we employ our developed framework to implement a scaled-down MobileNetV3 neural network for a classification application using the CIFAR-10 dataset, which comprises 32$\times$32 colour images. Following neural network computation, the CIFAR-10 dataset is categorized into ten distinct labels. The classification result determined by the memristor-based neural network is identified as the output with the highest current. The network weights are obtained from an offline server that was trained on the CIFAR-10 dataset. Table 1 compares the performance of our study with other neural networks utilizing novel computing paradigms. Our memristor-based MobileNetV3 neural network achieves a high classification accuracy of 90.36% on the CIFAR-10 dataset, comparable to traditional implementations using the PyTorch framework and exhibiting a significant improvement over previous works in this area.
IV-B Latency
Figure 7: (a) Latency and (b) power consumption of memristor-based MobileNetV3 neural network.
We refer to methods previously reported in the papers to analyze the latency and energy consumption of memristor-based neural networks [36, 37, 38, 19], to determine the latency of our proposed computing paradigm. In neural circuits based on memristors, the response time can be as quick as 100 ps [37]. However, the overall speed of the memristor-based circuit is constrained by the slew rate of the op-amps employed for converting current to voltage. So the latency of this network follows the following formula:
$T_{i}=(T_{m}+T_{o})*N_{m}+T_{r}$ (17)
where $T_{m}$ stands for the response time of the memristor crossbar, $T_{o}$ is the transition time of op-amps, $N_{m}$ denotes the number of memristor-based layers, $T_{r}$ represents the latency of other layers, and $T_{i}$ represents the latency of the inference process. The typical slew rate of the low-power op-amps is at 10 V/$\mu$s level [50]. $T_{r}$ includes the delays associated with modules such as the activation function layer, adders, and multipliers. We analyze the latency of these layers using information on existing devices. Therefore, the latency of this circuit can be as low as 1.24 $\mu$s, in comparison to 1.30 $\mu$s for the traditional dual op-amp solution. This is significantly faster than the traditional GPU approach, which has a latency of 0.1654 ms on a server equipped with an RTX 4090. As depicted in Figure 7(a), we also conducted inference time tests for single-image processing on a CPU (i7-12700), resulting in a latency of 3.3924 ms. To sum up, this computing paradigm achieves an extraordinary acceleration, boasting a speed-up of over 138 times compared to traditional GPU-based implementation and over 2827 times relative to conventional CPU-based solutions.
IV-C Energy Consumption
Figure 8 illustrates the distribution of memristor weights across different layers. As observed in Figure 8, the weights of the memristor crossbars predominantly range between -0.2 and 0.2. The power consumption of the memristor-based neural network is derived from the weights of the memristors. Since we have mapped the input data to ±2.5 mV, we can estimate the maximum predicted power consumption of a single memristor by assuming that all input data is at 2.5 mV and all weights are 0.2. Therefore, the maximum power consumption of memristors can be estimated at 1.1 $\mu$W. Consequently, we use the following formula to estimate the energy consumption:
$W_{i}=\sum U_{max}^{2}G_{max}*T_{m}+P_{o}*T_{o}+P{r}*T_{r}$ (18)
where $U_{max}$ denotes the maximum voltage across a single memristor,$G_{max}$ refers to the equivalent conductance of the memristors, $P_{o}$ represents the power consumption of op-amps, $P_{r}$ indicates the power used by other layers, and $W_{i}$ signifies the total power consumption. Based on the standard power consumption of devices, we estimate that a single complete forward inference of the neural network consumes 2.2 mJ. The comparative result of implementations using CPU, GPU, and dual op-amps is illustrated in Figure 7(b). As demonstrated in the figure, it achieves energy savings of 4.5 times compared to GPU implementation and 61.7 times compared to CPU implementation.
@@ IV-C Energy Consumption
V Conclusion
The memristor-based MobileNetV3 proposed in this paper demonstrates exceptional performance in terms of accuracy, latency, and energy consumption, underscoring the potential of this computing paradigm as a promising solution to overcome the computational resource constraints currently faced in machine learning. Additionally, the presentation of an automated framework for constructing and validating the memristor-based computing paradigm is poised to accelerate research in this field. Our results demonstrate that the approach we propose holds advantages not only compared to traditional CPU and GPU implementations but also against other novel computing paradigms, thus paving the way for the application of memristor-based computing paradigms in the realm of machine learning.